# Wake Vision: A Large-scale, Diverse Dataset and Benchmark Suite for TinyML Person Detection

### v1

@@ 3.3 Usability
Wake Vision is available through TensorFlow Datasets [33] and HuggingFace Datasets [22] to be easily accessed and used by the community. The images and labels are rehosted to ensure the dataset will not be changed over time due to dead links, making it a more viable benchmark. The rehosted labels are generated according to our default dataset configuration, further described in Appendix 0.F.
@@ 3.4 Comparison to Visual Wake Words
Wake Vision serves as a drop-in improvement over VWW, thanks to its size and our extensive filtering. To compare the benefits of Wake Vision against VWW, we train two identical models using the same recipe for the same number of steps, one utilizing VWW’s training set and the other employing Wake Vision’s training set. For this comparative evaluation, both trained models are MobileNetV2 architectures with a width modifier of 0.25, which have previously served as benchmark TinyML models [7]. After training these identical models using the respective datasets’ training sets with the same recipe for an equal number of steps, we cross-evaluate their performance on the corresponding test sets.
In Tab. 5 we demonstrate a direct 0.5% improvement on VWW’s own test set, which illustrates that Wake Vision is a direct improvement over VWW, not simply a domain shift. We additionally achieve a 2.41% improvement over VWW on Wake Vision’s Test set, indicating that the new test set is more challenging.
Table 5: Accuracy on the Wake Vision and VWW test sets by models trained on each dataset’s respective training set.
@@
VWW Wake Vision
Test VWW 88.00% 88.50%
Wake Vision 86.27% 88.68%
@@ 4 Fine Grain Benchmark Suite
Test sets generated directly from images posted to the internet are not reflective of real world use cases as the distribution is biased towards images people deems as worth sharing (e.g. well lit and framed). In contrast, person detection systems are deployed, and expected to work, in more challenging settings, such as low lighting. Due to this gap, a model that achieves high test accuracy may end up performing poorly in real world scenarios, after the design stage has ended.
To address this issue, we present a benchmark suite for TinyML person detection that tests a model’s robustness in challenging settings, enabling better analysis at the design stage. Additionally the benchmark suite measures the model’s performance across demographics, to ensure the model does not exhibit clear bias. The suite consists of five fine-grain benchmark sets, which we elaborate on below. Each of the five fine-grained benchmarks have been chosen based on a combination of its relevance to TinyML use cases and the availability of requisite metadata to generate the sets. Each benchmark set is a subset of the validation or test set filtered based on the criteria under test (e.g. persons far away from the camera).
The typical usage of the benchmark is to determine if a model is sufficiently accurate in the planned deployment setting. For example, a model designer may make different design choices for a use case where the subject is close to the camera and well lit, vs. the inverse setting. In this section, we describe the design of each individual benchmark and use the suite to compare a VWW model to a Wake Vision model. Tab. 6 reports the number of samples in each of the individual benchmark sets and the F1 scores of a Wake Vision and VWW model on each set.
4.1 Perceived Gender and Age
The Perceived Gender and Age fine grained datasets are generated from the Open Images More Inclusive Annotations for People (MIAP) extended fairness labels [32]. Underrepresented sub-groups typically constitute a small portion of a generic test set; therefore, top-line metrics often obscure a bias in a model until it is deployed [16]. This benchmark aims to evaluate a model separately on demographics that are underrepresented in the underlying dataset distribution, identifying bias. These labels are based on a perceived gender and age representation, and are not necessarily representative of a persons gender identity or age. Like the authors of the original MIAP labels, we do not support or condone the creation or deployment of gender or age specific classifiers based on this dataset.
Table 6: Wake Vision Fine Grain Benchmark Suite. We report the samples in each set as well as the F1 score of a Wake Vision model and a VWW model on the test set.
@@ 4.1 Perceived Gender and Age
Gender Female 675 2,152 0.94 0.89
Male 1,238 3,748 0.93 0.89
Unknown 1,718 5,231 0.78 0.79
Age Young 271 875 0.94 0.90
Middle 2,035 6,371 0.93 0.89
Older 91 271 0.96 0.90
Unknown 1,388 4,120 0.73 0.76
Distance Near 5,888 17,536 0.90 0.88
Medium 2,304 7,168 0.83 0.83
Far 1,152 3,328 0.39 0.49
Lighting Dark 3,036 8,872 0.88 0.83
Normal 12,992 39,327 0.88 0.86
Bright 936 3,083 0.85 0.85
Depictions Person 179 501 0.83 0.77
Non-Person 313 957 0.89 0.85
No Depiction 7,990 24,183 0.90 0.87
4.2 Distance
The distance fine grained datasets aim to test how the distance of persons in images impact the performance of ML systems. This benchmark is critical for correctly predicting a model’s deployed performance. If a person detection system is intended to recognize subjects at great distances, the system’s performance on the far away dataset will be more informative than its performance on the top-level test set. We create three datasets based on the percentage of the image the subject bounding box covers. The three sets are near (>60%), at a medium distance (10-60%), and far away (<10%).
4.3 Lighting
The lighting fine grained datasets aim to test the performance of ML systems across different lighting conditions. For instance, in scenarios such as surveillance cameras, outdoor robotics, or augmented reality applications, models need to be robust to varying lighting conditions, including low-light environments. As with distance, this benchmark provides insight into possible causes of significant performance degradation once deployed, such as low-light settings. We create three fine grained datasets of this type for dark, normal and bright lighting conditions respectively. All images in the original Wake Vision Vision validation and test set are classified into these datasets according to their lighting condition. We quantify lighting conditions by the average pixel values of images in greyscale, which is a simple but effective method for distinguishing lighting conditions [37]. We define low as an average pixel value less than 85, normal as between 85 and 170, and bright lighting conditions as greater than 170.
4.4 Depictions
@@ 4.5 Benchmark Results
Tab. 6 compares the results of a model trained on the Wake Vision (Quality) training set to an identical model trained on VWW. The Wake Vision model exhibits superior robustness across the challenging settings exercised by our benchmarking suite, further demonstrating the effectiveness of the dataset. For instance, on the “Depictions” benchmark, which evaluates performance on images containing persons, non-person objects, or no depictions, the Wake Vision model achieves an F1 score of 0.83 for person depictions, outperforming the VWW model’s 0.77. Similarly, for non-person depictions, the Wake Vision model scores 0.89 compared to the VWW model’s 0.85.
The Wake Vision model also showcases improved performance on the “Age” benchmark, with F1 scores of 0.94 for young subjects, 0.93 for middle-aged, and a notable 0.96 for older individuals, surpassing the VWW model’s scores of 0.90, 0.89, and 0.90, respectively. This highlights the Wake Vision dataset’s effectiveness in enhancing model robustness for detecting people across various age groups, particularly the elderly demographic. In addition to age and depictions, the Wake Vision model demonstrates resilience to challenging lighting conditions, achieving F1 scores of 0.88 for both dark and normal lighting scenarios, compared to the VWW model’s 0.83 and 0.86, respectively.
@@
5 Model Design Case Study
@@
5.1 Image Size vs. Model Width Scaling
@@
The results in Fig. 3 (left) show that when looking exclusively at the high-level metric (i.e., test accuracy), scaling the input image size has a similar impact as scaling the model size. However, as shown in Fig. 3 (right), when we consider only samples where the person is far away from the camera (i.e., the distance benchmark), we observe a much more significant impact when scaling the image size. In the case of distant subjects, the image size becomes the bottleneck.
@@ 5.2 Quantization
[…]
5.3 Grayscale
Converting a model’s input image channels to grayscale from RGB is a commonly employed optimization in the TinyML field [2] as it can substantially reduce a model’s memory consumption. We observed, however, that the grayscale optimization disproportionately impacts images on the brighter end of the spectrum as illustrated in Fig. 4. This further demonstrates the importance of fine grained analysis, as some real world deployment environments might be far brighter than the average Wake Vision test sample.
@@ 5.3 Grayscale
6 Ethical Considerations
The primary purpose of the Wake Vision dataset and benchmark suite is to mature the field of TinyML by giving researchers access to large-scale datasets and enabling fine-grain analysis. Beyond this purpose, Wake Vision can directly improve the quality of person detection systems that benefit society, such as saving energy by only turning on the lights in a room when a person is present. Additionally, TinyML person detection protects user privacy by running inference on-device and not transmitting private data over the network. However, it should be noted that person detection models can also be used for malicious purposes, such as weapons targeting systems or large-scale surveillance.
The images that make up Wake Vision are entirely sourced from Flickr. While we attempt to only use images under the consent given by the CC-BY 2.0 license (by sourcing images through Open Images[21, 19]), we make no representations or warranties regarding the license status of each image. It is possible that some of the millions of images have been uploaded to Flickr by a user without the right to distribute the images with a CC-BY 2.0 license.
Futhermore, the benchmark suite is intended to ensure the fairness and robustness of person detection models. While the goal of these benchmark suites is to build more robust and fair models, our demographic benchmarks could be misused to create systems to classify gender and age, potentially breaching personal privacy or perpetuating discrimination. We have attempted to mitigate this risk by not releasing training sets for the fine-grained benchmark datasets.
7 Conclusions
The field of TinyML research is currently bottlenecked by the lack of large, high-quality datasets. To alleviate this bottleneck we introduced Wake Vision—a large scale dataset and benchmark suite for person detection. Compared to VWW, the current state of the art dataset for person detection, Wake Vision is ~100x larger, easier to use, has a comparable amount of label errors, and is more reflective of real world use cases. These factors contribute to Wake Vision being a drop-in improvement over VWW. Fairness across demographics and robustness to challenging settings is essential for many person detection use cases. To facilitate evaluation of this, we released a Wake Vision benchmark suite of six fine-grained benchmark sets to test the fairness and robustness of person detection models. Lastly we demonstrate how the Wake Vision benchmark suite can be used to investigate the effects of different types of input and model scaling and highlight the importance of fine grain analysis when designing an model.
Appendix 0.A Anonymized Repo
The code used for the paper is available at this anonymous repository: https://anonymous.4open.science/r/Wake_Vision-0D15/README.md
Appendix 0.B Flowchart of Bounding Box Filtering Process
Appendix 0.C Fine Grained Benchmark Images
@@
Appendix 0.D Open Image Label Distribution
@@
Appendix 0.E Dataset Access and Organization
Wake Vision is available directly online, via TFDS, or Hugging Face Datasets. In order to maintain author anonymity, we do not link to the dataset.
The dataset is organized as a set of compressed tar files containing images, and a series of label CSVs. The label CSVs are organized such that the file name of the image is the identifying index. The person label is provided as is a flag to indicate if the image is of a depiction. The Validation and Test label CSVs also have flags that denote a sample’s inclusion into a fine-grain benchmark set (e.g. Distance-Near). This structure makes it easy to access just the required data without requiring a full download. It also ensures the dataset can be easily updated as new versions are introduced.
Appendix 0.F Label Generation Details
Person Labels. The most straightforward Open Images label classes to label as person in Wake Vision are the "person" label and its subcategories (listed in Appendix 0.I). All of these are relabelled as persons in Wake Vision. These labels are present as both image-level labels and bounding box labels.

### v2

@@ 3.3 Usability
Wake Vision is available through TensorFlow Datasets [28] and HuggingFace Datasets [17] to enable easy access and use by the community. The images and labels are rehosted to ensure the dataset will not be changed over time due to dead links, making it a more stable benchmark. The rehosted labels are generated according to our default dataset configuration, further described in Sec. I.
@@ 3.4 Comparison to Visual Wake Words
Wake Vision serves as a drop-in improvement over VWW, thanks to its size and our extensive filtering. To compare the benefits of Wake Vision against VWW, we train two identical MobileNetv2 [26] models with a width modifier of 0.25 for 200,000 steps on 224x224x3 images using AdamW [20]. We use a lr of 0.002 with a cosine decay and a weight decay of 4E-6. One model uses VWW’s training set, and the other employs Wake Vision’s Quality training set. After training these identical models using the respective datasets’ training sets with the same recipe for an equal number of steps, we cross-evaluate their performance on the corresponding test sets.
Table 4 shows a 0.26% improvement on VWW’s own test set, which illustrates that Wake Vision is a direct improvement over VWW, not simply a domain shift. We additionally achieve a 1.1% improvement over VWW on Wake Vision’s Test set, indicating that the new test set is more challenging. Table 3 also shows the performance of a the Wake Vision (Combined) model described in section 3.1. This model achieves an average 0.75% improvement over the Wake Vision (Quality) trained model on the VWW test set and an average 0.83% improvement on the Wake Vision test set. The Wake Vision (Combined) model is trained for longer than the VWW and Wake Vision (Quality) models, but shows the value of using the training sets in unison even on the VWW test set.
Table 4: Accuracy on the Wake Vision and VWW test sets by models trained on the VWW, Wake Vision (Quality) and Wake Vision (Combined) training sets.
@@
VWW Wake Vision (Quality) Wake Vision (Combined)
Test VWW 88.33$\pm$0.29 88.59$\pm$0.1789.34$\pm$0.02%
Wake Vision 83.79$\pm$0.2384.89$\pm$0.1185.72$\pm$0.04%
3.5 Dataset Quality Vs. Size
We investigate the impact of dataset quality and size of Resnet [10] style models of varying capacity. We measure dataset quality by the approximate rate of label errors. The Wake Vision (Quality) training set has an estimated error rate of around 7%. We simulate higher error rate datasets (15% and 30%) by flipping the binary label with a certain probability. Appendix D gives the derivation for the flip probability. We also sweep the dataset size by taking a slice of Wake Vision (Quality). Each model is trained for 50,000 steps on 224x224x3 images using AdamW [20], a learning rate of 0.001 with a cosine schedule and a weight decay of 0.004. Figure 3 illustrates the experiment’s results. We observe that smaller models (leftmost figure) are more sensitive to higher error rates than large models. The smallest model’s accuracy drop is 1.3%, going from the base train (quality) error rate of around 7% to 15% compared to only a 0.5% accuracy loss for the largest model. In contrast, large models benefit more from big datasets and start to over fit on the smaller slices of the training data.
Figure 3: Comparing the impact of dataset error rate and size on ResNet models of varying capacity. The $y$-axis is the Wake Vision test accuracy and the $x$-axis is the percentage of Wake Vision Train (Quality) used in training. The error % indicates the expected rate of incorrectly labeled samples.
@@ 4 Fine Grain Benchmark Suite
Test sets generated directly from images posted to the internet are not reflective of real world use cases as the distribution is biased towards images people deems as worth sharing (e.g. well lit and framed). In contrast, person detection systems are deployed, and expected to work, in more challenging settings, such as low lighting. Due to this gap, a model that achieves high test accuracy may perform poorly in real-world scenarios after the design stage has ended. To address this issue, we present a benchmark suite for person detection that tests a model’s robustness in challenging settings, enabling better analysis at the design stage. Additionally, the suite measures the model’s performance across demographics to ensure it does not exhibit clear bias. The suite consists of five fine-grained benchmark sets. Each of the five fine-grained benchmarks have been chosen based on a combination of its relevance to TinyML use cases and the availability of requisite metadata to generate the sets. Each benchmark set is a subset of the respective validation or test set filtered based on the criteria under test. The benchmark determines whether a model is sufficiently accurate in the planned deployment setting. For example, a model designer may make different design choices for a use case where the subject is close to the camera and well-lit compared to the inverse setting. Appendix C contains a case study showing how the fine grained evaluation sets can be used to design robust models. Table 5 reports the number of samples in each of the individual benchmark sets and the F1 scores of a Wake Vision and VWW model on each set.
Perceived Gender and Age
These fine grained datasets are generated from the Open Images More Inclusive Annotations for People (MIAP) extended fairness labels [27]. Underrepresented sub-groups typically constitute a small portion of a generic test set; therefore, top-line metrics often obscure a bias in a model until it is deployed [12]. This benchmark aims to evaluate a model separately on demographics that are underrepresented in the underlying dataset distribution, identifying bias. These labels are based on a perceived gender and age representation and are not necessarily representative.
Table 5: Wake Vision Fine Grain Benchmark Suite. We report the samples in each set and the average F1 score across three Wake Vision models and three VWW models on the benchmark set.
@@
Gender Female 684 2,157 0.93 0.89
Male 1,310 3,918 0.91 0.88
Unknown 1,612 4,940 0.77 0.78
Age Young 275 884 0.94 0.90
Middle 2,133 6,595 0.91 0.88
Older 90 276 0.94 0.89
Unknown 1,299 3,837 0.71 0.74
Distance Near 5,457 16,333 0.91 0.89
Medium 2,213 6,876 0.85 0.84
Far 398 1,140 0.59 0.67
Lighting Dark 3,255 9,420 0.85 0.81
Normal 14,315 43,010 0.82 0.82
Bright 1,012 3,332 0.82 0.82
Depictions Person 356 978 0.71 0.66
Non-Person 352 1101 0.86 0.82
No Depiction 8,583 25,802 0.87 0.85
Distance
The distance datasets test how the distance of persons in images impact the performance of ML systems. This benchmark is critical for correctly predicting a model’s deployed performance. If a person detection system is intended to recognize subjects at great distances, the system’s performance on the far away dataset will be more informative than its performance on the top-level test set. We create three datasets based on the percentage of the image the subject bounding box covers. The three sets are near (>60%), at a medium distance (10-60%), and far away (<10%).
Lighting
The lighting datasets aim to test the performance of ML systems across different lighting conditions. In scenarios like security cameras, outdoor robotics, or augmented reality applications, models must be robust to varying lighting conditions, including low-light environments. We create three fine grained datasets of this type for dark, normal, and bright lighting conditions, respectively. We quantify lighting conditions by the average pixel values of images in greyscale, which is a simple but effective method for distinguishing lighting conditions [33]. We define low as an average pixel value less than 85, normal as between 85 and 170, and bright lighting conditions as greater than 170.
Depictions
@@ 4.1 Benchmark Results
Table 5 compares the results of a model trained on the Wake Vision (Quality) training set to an identical model trained on VWW. The Wake Vision model exhibits superior robustness across the challenging settings exercised by our benchmarking suite, further demonstrating the effectiveness of the dataset. For instance, on the “Depictions” benchmark, which evaluates performance on images containing persons, non-person objects, or no depictions, the Wake Vision model achieves an F1 score of 0.71 for person depictions, outperforming the VWW model’s 0.66. Similarly, the Wake Vision model scores 0.86 for non-person depictions compared to the VWW model’s 0.82. The Wake Vision model also showcases improved performance on the “Age” benchmark, with F1 scores of 0.94, 0.91, and 0.94 for young, middle, and older individuals, surpassing the VWW model’s scores of 0.90, 0.88, and 0.89, respectively. This highlights the dataset’s effectiveness in enhancing model robustness for detecting people across various age groups, particularly the elderly demographic. In addition, the Wake Vision model demonstrates resilience to challenging lighting conditions, achieving F1 scores of 0.85 in dark lighting scenarios, compared to the VWW’s 0.81.
@@
5 Ethical Considerations
The Wake Vision dataset and benchmark suite aim to advance TinyML research by providing large-scale datasets and enabling fine-grained analysis of TinyML systems. Wake Vision can directly improve the quality of person detection systems that benefit society, such as energy-saving applications, while protecting user privacy through on-device inference. However, person detection models can also be used for malicious purposes, such as weapons targeting or large-scale surveillance.
The images in Wake Vision are sourced from Flickr through Open Images[16, 14] under the CC-BY 2.0 license. While we have made efforts to ensure the images are properly licensed, we cannot guarantee the license status of each image. It is possible that some images have been uploaded without the right to distribute under the CC-BY 2.0 license. Given sufficient resources, individuals in the dataset could potentially be identified based on biometric data. The dataset may also contain offensive content, as manually verifying millions of images is infeasible.
The benchmark suite aims to ensure the fairness and robustness of person detection models. However, the demographic benchmarks could potentially be misused to create systems that classify gender and age, breaching personal privacy or perpetuating discrimination. To mitigate this risk, we have not released training sets for the fine-grained benchmark datasets.
6 Conclusions
The lack of large, high-quality datasets has been a significant bottleneck in the field of TinyML research. To address this issue, we introduced Wake Vision, a large-scale dataset and benchmark suite for person detection that surpasses the current state-of-the-art dataset, VWW, in terms of size (~100x larger), ease of use, label accuracy, and real-world relevance. These improvements make Wake Vision a drop-in replacement for VWW. Recognizing the importance of fairness across demographics and robustness to challenging settings in person detection use cases, we also released a Wake Vision benchmark suite comprising six fine-grained benchmark sets to evaluate the fairness and robustness of person detection models. Furthermore, our investigation into the trade-off between dataset quality and quantity for TinyML systems suggests that dataset quality is likely more crucial for TinyML systems compared to conventional ML systems. The Wake Vision dataset and benchmark suite aim to facilitate the development of more accurate, fair, and robust person detection models in TinyML.
7 Acknowledgements
This work was partially supported by the Google TPU Research Cloud program. We would like to thank Andrew Howard of Google Deepmind for his input and support in this effort.
Appendix A Code Repository
The code used to generate the Wake Vision dataset is available at the following GitHub repository: https://github.com/harvard-edge/Wake_Vision
Appendix B Flowchart of Bounding Box Filtering Process
@@
Appendix C Model Design Case Study
@@
C.1 Image Size vs. Model Width Scaling
@@
The results in Table 5 (left) show that when looking exclusively at the high-level metric (i.e., test accuracy), scaling the input image size has a similar impact as scaling the model size. However, as shown in Figure 5 (right), when we consider only samples where the person is far away from the camera (i.e., the distance benchmark), we observe a much more significant impact when scaling the image size. In the case of distant subjects, the image size becomes the bottleneck.
@@ C.2 Quantization
Our findings show negligible degradation in performance across all benchmarks (±$\pm 0.004$ F1) when employing int8 quantization, even on outlier sets. This result contradicts the previously observed disproportionate impact of quantization on underrepresented subsets. We speculate that person detection may be a relatively simple task, potentially explaining why we do not observe this specific property of quantization in our experiments. Given the negligible performance degradation and the substantial latency, memory, and model size benefits of quantization, we conclude that quantization is a win for person detection.
C.3 Grayscale
Converting a model’s input image channels to grayscale from RGB is a commonly employed optimization in the TinyML field [2] as it can substantially reduce a model’s memory consumption. We observed, however, that the grayscale optimization disproportionately impacts images on the brighter end of the spectrum as illustrated in Figure 6. This further demonstrates the importance of fine grained analysis, as some real world deployment environments might be far brighter than the average Wake Vision test sample.
@@ C.3 Grayscale
Appendix D Inducing Label Errors
The goal is to make a single pass through the dataset and flip the labels of a binary classification dataset such that the expected label error rate is $d$. There is an underlying rate of label errors $e$. If we flip one of these underlying errors, we correct it, thereby inadvertently decreasing the label error rate. After flipping labels with a probability of $p$ we can claim that the likelihood of a single label being correct is the probability that we flipped the label and it was originally an error plus the probability that we didn’t flip the label and it was originally correct: $1-d=p*e+(1-p)(1-e)$. Then solving for $p$ gives $p=(e-d)/(2e-1)$. A current flaw of this method is that the injected label errors are not consistent between epochs, which would likely be less destructive to a model’s accuracy since the same errors are not reinforced each epoch. This could also potentially explain why large models in Fig. 3 don’t overfit on training data with higher label errors, as the inconsistent label noise has a regularizing effect.
Appendix E Fine Grained Benchmark Images
@@
Appendix F Open Image Label Distribution
@@
Appendix G Manual Labeling
The validation and test sets were manually labeled through the crowd-sourced labeling platform Scale Rapid(https://scale.com/. Figure 8 shows a screenshot of the labeling menu. The labelers were instructed to label an image as a "person" if a person was present anywhere in the image and "no person" if no visible person was present. The labelers also indicated if the image contained a depiction of a person, which meant that sample was labeled "no person" in Wake Vision, or a "picture of a picture/reflection", which we used for metadata.
The cost per image was \$0.10 USD and is set based on Scale’s pricing structure. Each image is labeled by 3 different labelers to form a consensus. The total cost of labeling was \$7089.8 USD. The authors do not know the hourly rate paid to each labeler, but Scale lists \$18/hr on job postings at the time of writing. For context on the difficulty of the task, the authors were able to average around 500 images per hour at a reasonable pace when estimating error rates.
Figure 8: A screenshot of the labeling menu used to manually label the validation and test sets.
Appendix H Dataset Access and Organization
Wake Vision is available online at the Harvard Dataverse, via TFDS, or Hugging Face Datasets. More information can be found on the https://wakevision.ai webpage.
The dataset is organized as a set of compressed tar files containing images, and a series of label CSVs. The label CSVs are organized such that the file name of the image is the identifying index. CSVs for all dataset splits include the person and depiction label. The Validation and Test label CSVs also have flags that denote a sample’s inclusion into a fine-grain benchmark set (e.g. Distance-Near). This structure makes it easy to access just the required data without requiring a full download. It also ensures the dataset can be easily updated as new versions are introduced.
Appendix I Label Generation Details
Person Labels. The most straightforward Open Images label classes to label as person in Wake Vision are the "person" label and its subcategories (listed in appendix L). All of these are relabelled as persons in Wake Vision. These labels are present as both image-level labels and bounding box labels.